OpenStudy (anonymous):

Determine whether the sequence converges or diverges. If it converges, give the limit 11, 22, 44, 88, ...

OpenStudy (anonymous):

they get bigger and bigger right?

OpenStudy (victoriasushchik):

the common ratio is r = 2... so the sequence is divergent...

OpenStudy (anonymous):

how do i find the limit now

OpenStudy (victoriasushchik):

convergence the common ratio is -1 < r < 1 so if thats the case the limiting sum is S∞=a1−r

OpenStudy (anonymous):

\[a _{n}=11*2^{n-1}\] \[\lim_{n \rightarrow \infty }a _{n}\rightarrow \infty and \not~\rightarrow 0\] so series diverges

OpenStudy (idku):

For sequence, actually -1 < r ≤ 1

OpenStudy (idku):

it will not converge for r=-1, because it will always alternate. and if r=1, then the sequence (not series) converges.

OpenStudy (idku):

(this is where the ratio test comes from, to show that series behaves in a way that its "r" satisfies |r|<1, but this is for series, sequence is good when r=1)

OpenStudy (anonymous):

@idku I'm still not understanding so the ratio for mine is 2 right?

OpenStudy (idku):

Look at this way (as satelite offered). YOur terms will get bigger and bigger (because you are mutliplying times 2 every time). And this will get larger and larger with no limit...

OpenStudy (idku):

So, does your sequence tend to some (particular) value, or does the sequence not?

OpenStudy (idku):

(if the sequence becomes infinitely large, that means it will diverge)

OpenStudy (anonymous):

yeah it seems like it will go forever @idku